OpenStudy (anonymous):

TRUE or FALSE! :) tan^2 x + sec^2 x = 1 for all values of x. ***My answer: False is that right? @terenzreignz :)

terenzreignz (terenzreignz):

that's right :) -tan^2 x + sec^2 x = 1 is true for all values. :)

OpenStudy (anonymous):

ohhh i see :) so this is TRUE right?

terenzreignz (terenzreignz):

No, I said you were correct :P

terenzreignz (terenzreignz):

I said it would be true for all possible values of x if it were -tan^2 x instead

OpenStudy (anonymous):

TRUE or FALSE! :) cos2a=1-2sin^2 a for all values of a. ***MY answer: FALSE is that right? :) and ohhh okay haha i see :) yay! thanks! :D

OpenStudy (anonymous):

what do u think of that one @terenzreignz :)

OpenStudy (anonymous):



terenzreignz (terenzreignz):

Well, problem. You know the pythagorean identity \[\large \cos^2x + \sin^2 x = 1\] from which you can derive \[\large \color{red}{\cos^2x }= \color{blue}{1-\sin^2x}\] And the eerily similar formula for the double angle identity of cosine \[\large \cos(2x) = \color{red}{\cos^2x }- \sin^2x\] Here you can replace the \(\large \cos^2x\) as prescribed by our identity... \[\large \cos(2x) = \color{blue}{1-\sin^2x}-\sin^2x\]\[\large \cos(2x) = 1-2\sin^2x\]

OpenStudy (anonymous):

ohhh okay... so it's true then right ? cuz we end up with this? cos2a=1-2sin^2a just a instead of the x tho? so TRUE?!

terenzreignz (terenzreignz):

Yes, true.

OpenStudy (anonymous):

okay! thanks! :D